I need help with Linear Equation with two variables
A rectangular lot whose perimeter is 480 feet is fenced along three sides. An expensive fencing along the lotʹs
length costs $16 per foot , and an inexpensive fencing along the two side widths costs only $9 per foot. The total
cost of the fencing along the three sides comes to $4040. What are the lotʹs dimensions?

Answer by stanbon(75887)

A rectangular lot whose perimeter is 480 feet is fenced along three sides. An expensive fencing along the lotʹs length costs $16 per foot , and an inexpensive fencing along the two side widths costs only $9 per foot. The total
cost of the fencing along the three sides comes to $4040. What are the lotʹs dimensions?
---
Let i be the inexpensive and e the expensive.
-------
Measure Equation: e + 2i = 480 ft.
Value Equation:: 16e + 9i = 4040
-------------------------------------------
Multiply thru the Measure Eq. by 16 to get:
16e + 32i = 16*480
-----
Subtract the Value Eq. from that and solve for "i":
23i = 3640
i = 158.26 ft (length of the inexpensive dimension)
----
Substitute into e + 2i = 480 to solve for "e":
e + 2*158.26 = 480
e = 163.48 ft (length of the expensive dimension)
===================================================
Cheers,
Stan H.
========